OpenStudy (anonymous):

write a formula for f of g of h, where f(x) = x-2, g(x)=-4x, and h(x)=4-x

OpenStudy (anonymous):

Is it f(g(h(x))? Then, first, find a formula for g(h(x)). Sub h in g for x to get: g(h(x)) = -4(4-x) Then sub this in f(x) to get: f(g(h(x)) = -4(4-x) - 2 You can simplify it now

OpenStudy (anonymous):

f(g(h(x)))=f(g(4-x))=f(-4(4-x))=f(-16+4x)=(-16+4x)-2=4x-18. When you write f(x)=y, this means that f maps x to y. In this case, x is replaced by some other function.